The sum of two numbers is 33. Their difference is 11. What are the two numbers?
Answer by praseenakos@yahoo.com(507)

QUESTION:

The sum of two numbers is 33. Their difference is 11. What are the two numbers?

ANSWER:

Assume that two mumbers are x and y and suppose x is the bigger number.


It is given that their sum is 33.

==> x + y = 33 -----------------(1)


Their difference is 11

==> x- y = 11 -------------------(2)


Add (1) and (2)


==> x + y + x - y = 33 + 11


==> x + x + y - y = 44


==> 2x = 44


Divide both sides of the expression by 2.


==> 2x/2 = 44/2


==> x = 22


So the bigger number is 22

Given that sum of these numbers = 33

So the other number is = 33- 22 = 11


So the numbers are 22 and 11.


Hope you understood.

Regards.

Praseena.